# presep= ]e1,e2,e3

### v1

@@ A.4 Proof of Theorem 4 (Functional CLT)
$\frac{r_{n}}{\varepsilon\sqrt{k_{n}h_{n}}}\Biggl{(}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)^{3}1_{\left\{Y_{0}>su_{n}\right\}}\right]$
$\qquad\qquad+\sum_{j=1}^{r_{n}}\left(1-\frac{j}{r_{n}}\right)\biggr{\{}\mathbb{E}\left[K\left(\frac{x-X_{j}}{h_{n}}\right)^{2}K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}1_{\left\{Y_{0}>su_{n}\right\}}\right]$
$\qquad\qquad+\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)^{2}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}}1_{\left\{Y_{j}>su_{n}\right\}}\right]\biggr{\}}\Biggr{)},$
where the first term is order $o\left(r_{n}h_{n}\overline{F}^{x}(u_{n})\right)$ as above using condition 1 and Lemma 5. The leading term in the sum is dealt with as in the proof of Lemma 6, from which we infer that the expression is $o\left(r_{n}h_{n}\overline{F}^{x}(u_{n})\right)$ and consequently $I_{1}=o\left(r_{n}h_{n}\overline{F}^{x}(u_{n})\right)$.
@@
$\sum_{j=m}^{r_{n}}\frac{\mathbb{E}\left[K\left(\frac{x-X_{j}}{h_{n}}\right)K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}1_{\left\{Y_{0}>su_{n}\right\}}\right]}{h_{n}\overline{F}^{x}(u_{n})}\leq\delta.$
@@
$\sum_{j=1}^{m}\left(m-j\right)\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}}1_{\left\{Y_{j}>su_{n}\right\}}1_{\left\{\lvert\overline{S}_{n,1}^{x}(s)\rvert>\varepsilon\sqrt{k_{n}h_{n}}\right\}}\right]$
$\leq m\sum_{j=1}^{m}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}}1_{\left\{Y_{j}>su_{n}\right\}}\right]$
$\leq\frac{m}{r_{n}}r_{n}h_{n}\overline{F}^{x}(u_{n})\sum_{j=1}^{r_{n}}\frac{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}}1_{\left\{Y_{j}>su_{n}\right\}}\right]}{h_{n}\overline{F}^{x}(u_{n})}$
$=\frac{\mathcal{O}\left({r_{n}h_{n}\overline{F}^{x}(u_{n})}\right)}{r_{n}}=o\left(r_{n}h_{n}\overline{F}^{x}(u_{n})\right).$
@@
$\sum_{i=1}^{r_{n}}\sum_{j=i+m}^{r_{n}}\mathbb{E}\left[K\left(\frac{x-X_{i}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{i}>su_{n}\right\}}1_{\left\{Y_{j}>su_{n}\right\}}1_{\left\{\lvert\overline{S}_{n,1}^{x}(s)\rvert>\varepsilon\sqrt{k_{n}h_{n}}\right\}}\right]\leq\delta r_{n}h_{n}\overline{F}^{x}(u_{n}).$
@@
$\limsup_{n\to\infty}\frac{I_{2}}{r_{n}h_{n}\overline{F}^{x}(u_{n})}\leq\delta,$
and since $\delta$ can be made arbitrarily small we conclude that $I_{2}=o\left(r_{n}h_{n}\overline{F}^{x}(u_{n})\right)$ which concludes the proof. ∎
Now we show that terms constructed from the incomplete block $\left\{\left(X_{m_{n}r_{n}+1},Y_{m_{n}r_{n}+1}\right),\ldots,\left(X_{n},Y_{n}\right)\right\}$ are asymptotically negligible.
@@
$\limsup_{n\to\infty}\sup_{s\in[a,b]}\left\lvert\sum_{j=m_{n}r_{n}+1}^{n}\frac{K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}-\mathbb{E}\left[K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}\right]}{\sqrt{nh_{n}\overline{F}^{x}(u_{n})}}\right\rvert=0$
@@
Since $s\mapsto K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}$ is monotonically decreasing, and $(n-m_{n}r_{n})\leq r_{n}$, i.e. the last incomplete block consists of fewer terms than a complete block, we get
$\sup_{s\in[a,b]}\left\lvert\sum_{j=m_{n}r_{n}+1}^{n}\frac{K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}-\mathbb{E}\left[K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}\right]}{\sqrt{nh_{n}\overline{F}^{x}(u_{n})}}\right\rvert$
$\leq\frac{1}{\sqrt{nh_{n}\overline{F}^{x}(u_{n})}}\Bigg{\{}\sum_{j=m_{n}r_{n}+1}^{n}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>au_{n}\right\}}+r_{n}\mathbb{E}\left[K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>au_{n}\right\}}\right]\Bigg{\}}.$
@@
$\frac{1}{nh_{n}\overline{F}^{x}(u_{n})}\operatorname{var}\left(\sum_{j=1}^{(n-m_{n}r_{n})}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>au_{n}\right\}}\right)$
$\leq\frac{1}{nh_{n}\overline{F}^{x}(u_{n})}\operatorname{var}\left(\sum_{j=1}^{r_{n}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>au_{n}\right\}}\right)=\frac{1}{m_{n}}\operatorname{var}\left(\widetilde{\mathbb{V}}^{{\dagger},x}_{n}(a)\right)\to 0,$
where we used that the joint process $\left\{\left(X_{j},Y_{j}\right)\right\}$ is stationary, Lemma 5 and Lemma 6 . Further, by condition 1
$\frac{r_{n}}{\sqrt{nh_{n}\overline{F}^{x}(u_{n})}}\mathbb{E}\left[K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>au_{n}\right\}}\right]=\frac{r_{n}}{\sqrt{nh_{n}\overline{F}^{x}(u_{n})}}\mathcal{O}\left(h_{n}\overline{F}^{x}(u_{n})\right)=\frac{\mathcal{O}\left(r_{n}h_{n}\overline{F}^{x}(u_{n})\right)}{\sqrt{nh_{n}\overline{F}^{x}(u_{n})}}\to 0.$
@@
$\sum_{j=1}^{r_{n}}\frac{K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}}{\sqrt{nh_{n}\overline{F}^{x}(u_{n})}}.$
@@
$m_{n}\mathbb{E}\left[\left(\sum_{j=1}^{\ell_{n}}\frac{K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}}{\sqrt{nh_{n}\overline{F}^{x}(u_{n})}}\right)^{2}\right]\leq\frac{l_{n}}{r_{n}}\operatorname{var}\left(\widetilde{\mathbb{V}}^{{\dagger},x}_{n}(s)\right)\to 0.$
@@
Assume conditions 1-4, 1-3. Then $\widetilde{\mathbb{V}}^{x}_{n}$ converges weakly to the centred Gaussian process ${\mathbb{V}}^{x}$ in $\mathbb{D}((0,\infty))$ endowed with the $J_{1}$-topology, and the process $\mathbb{V}^{x}$ admits an almost surely continuous version.
@@
$\widetilde{\mathbb{V}}^{x}_{n,1}=\sum_{\begin{subarray}{c}1\leq i\leq m_{n}\\ i\ \text{odd}\end{subarray}}\frac{\overline{S}_{n,i}^{x}}{\sqrt{nh_{n}\overline{F}^{x}(u_{n})}},\quad\ \ \widetilde{\mathbb{V}}^{x}_{n,2}=\sum_{\begin{subarray}{c}1\leq i\leq m_{n}\\ i\ \text{even}\end{subarray}}\frac{\overline{S}_{n,i}^{x}}{\sqrt{nh_{n}\overline{F}^{x}(u_{n})}}$
$\widetilde{\mathbb{V}}^{{\dagger},x}_{n,1}=\sum_{\begin{subarray}{c}1\leq i\leq m_{n}\\ i\ \text{odd}\end{subarray}}\frac{\overline{S}_{n,i}^{{\dagger},x}}{\sqrt{nh_{n}\overline{F}^{x}(u_{n})}},\quad\widetilde{\mathbb{V}}^{{\dagger},x}_{n,2}=\sum_{\begin{subarray}{c}1\leq i\leq m_{n}\\ i\ \text{even}\end{subarray}}\frac{\overline{S}_{n,i}^{{\dagger},x}}{\sqrt{nh_{n}\overline{F}^{x}(u_{n})}}$
@@
$\widetilde{\mathbb{V}}^{x,\text{incomplete}}_{n}(s):=\sum_{j=m_{n}r_{n}+1}^{n}\frac{K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}-\mathbb{E}\left[K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}\right]}{\sqrt{nh_{n}\overline{F}^{x}(u_{n})}}.$
@@
The summands of $\widetilde{\mathbb{V}}^{{\dagger},x}_{n}$ are i.i.d processes indexed on $[a,b]$ with càdlàg sample paths, the latter implying implying separability of each summand. Consequently we can apply Theorem 19 to show asymptotic equicontinuity. The Lindeberg condition was already shown in Theorem 10. With $v_{n}=nh_{n}\overline{F}^{x}(u_{n})$ define the random metric $\mathrm{d}_{n}$ on $[a,b]$ by
@@
$\mathbb{E}\left[\mathrm{d}_{n}^{2}(s,t)\right]=\frac{m_{n}}{nh_{n}\overline{F}^{x}(u_{n})}\mathbb{E}\left[\left\{\overline{S}_{n,i}^{{\dagger},x}(s)-\overline{S}_{n,i}^{{\dagger},x}(t)\right\}^{2}\right]$
$\leq\frac{m_{n}}{nh_{n}\overline{F}^{x}(u_{n})}\left(\mathbb{E}\left[\overline{S}_{n,i}^{{\dagger},x}(s)^{2}\right]-\mathbb{E}\left[\overline{S}_{n,i}^{{\dagger},x}(t)^{2}\right]\right)\to\left(s^{-\alpha_{x}}-t^{-\alpha_{x}}\right)g(x)\int K^{2}\left(u\right)\mathrm{d}u,$
@@
$\lim_{n\to\infty}\sup_{\begin{subarray}{c}s,t\in[a,b]\\ \lvert s-t\rvert<\delta_{n}\end{subarray}}\mathbb{E}\left[\mathrm{d}_{n}^{2}(s,t)\right]\leq\lim_{n\to\infty}\sup_{\begin{subarray}{c}s,t\in[a,b]\\ \lvert s-t\rvert<\delta_{n}\end{subarray}}\left\{\frac{m_{n}}{nh_{n}\overline{F}^{x}(u_{n})}\left(\mathbb{E}\left[\overline{S}_{n,i}^{{\dagger},x}(s)^{2}\right]-\mathbb{E}\left[\overline{S}_{n,i}^{{\dagger},x}(t)^{2}\right]\right)\right\}=0.$
@@
$\mathbb{Z}_{n}\left(\mathcal{E}_{s}\right)$ $=v_{n}^{-1}\sum_{i=1}^{m_{n}}\mathcal{E}_{s}\left(u_{n}^{-1}\boldsymbol{Y}^{{\dagger}}_{(i-1)r_{n}+1,ir_{n}},h_{n}^{-1}\left(\boldsymbol{x}-\boldsymbol{X}^{{\dagger}}_{(i-1)r_{n}+1,ir_{n}}\right)\right)=\frac{1}{nh_{n}\overline{F}^{x}(u_{n})}\sum_{i=1}^{m_{n}}S_{n,i}^{{\dagger},x}(s),$
@@
$\widetilde{d}_{n}^{2}\left(\mathcal{E}_{s},\mathcal{E}_{t}\right)$ $=v_{n}^{-1}\sum_{i=1}^{m_{n}}\Biggl{\{}\mathcal{E}_{s}\left(u_{n}^{-1}\boldsymbol{Y}^{{\dagger}}_{(i-1)r_{n}+1,ir_{n}},h_{n}^{-1}\left(\boldsymbol{x}-\boldsymbol{X}^{{\dagger}}_{(i-1)r_{n}+1,ir_{n}}\right)\right)$
$\qquad\qquad\quad-\mathcal{E}_{t}\left(u_{n}^{-1}\boldsymbol{Y}^{{\dagger}}_{(i-1)r_{n}+1,ir_{n}},h_{n}^{-1}\left(\boldsymbol{x}-\boldsymbol{X}^{{\dagger}}_{(i-1)r_{n}+1,ir_{n}}\right)\right)\Biggr{\}}^{2},$
@@
$\boldsymbol{G}\left(\boldsymbol{y},\boldsymbol{x}\right)=\sup_{s,t\in[a,b]}\left\lvert\mathcal{E}_{s}\left(\boldsymbol{y},\boldsymbol{x}\right)\right\rvert=\sup_{s,t\in[a,b]}\left\lvert\sum_{j\in\mathbb{Z}}1_{\left\{y_{j}>s\right\}}K(x_{j})\right\rvert=\sum_{j\in\mathbb{Z}}1_{\left\{y_{j}>a\right\}}K(x_{j}).$
@@
$\limsup_{n\to\infty}\frac{m_{n}}{nh_{n}\overline{F}^{x}(u_{n})}\mathbb{E}\left[\boldsymbol{G}^{2}\left(u_{n}^{-1}\boldsymbol{Y}^{{\dagger}}_{1,r_{n}},h_{n}^{-1}\left(\boldsymbol{x}-\boldsymbol{X}^{{\dagger}}_{1,r_{n}}\right)\right)\right]$
$=\limsup_{n\to\infty}\frac{m_{n}}{nh_{n}\overline{F}^{x}(u_{n})}\mathbb{E}\left[\left(\sum_{j=1}^{r_{n}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>au_{n}\right\}}\right)^{2}\right]$
$=\limsup_{n\to\infty}\operatorname{var}\left(\widetilde{\mathbb{V}}^{{\dagger},x}_{n}(a)\right)=a^{-\alpha_{x}}g(x)\int K^{2}(u)\mathrm{d}u<\infty.$

### v2

@@ A.4 Proof of Theorem 4 (Functional CLT)
$\frac{r_{n}}{\varepsilon\sqrt{k_{n}h_{n}}}\Biggl(\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)^{3}1_{\left\{Y_{0}>su_{n}^{x}\right\}}\right]$
$\qquad\qquad+\sum_{j=1}^{r_{n}}\left(1-\frac{j}{r_{n}}\right)\biggr\{\mathbb{E}\left[K\left(\frac{x-X_{j}}{h_{n}}\right)^{2}K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}1_{\left\{Y_{0}>su_{n}^{x}\right\}}\right]$
$\qquad\qquad+\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)^{2}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}}1_{\left\{Y_{j}>su_{n}^{x}\right\}}\right]\biggr\}\Biggr),$
where the first term is order $o(r_{n}h_{n}\overline{F}^{x}(u_{n}^{x}))$ as above using condition 1 and Lemma 5. The leading term in the sum is dealt with as in the proof of Lemma 6, from which we infer that the expression is $o(r_{n}h_{n}\overline{F}^{x}(u_{n}^{x}))$ and consequently $I_{1}=o(r_{n}h_{n}\overline{F}^{x}(u_{n}^{x}))$.
@@
$\sum_{j=m}^{r_{n}}\frac{\mathbb{E}\left[K\left(\frac{x-X_{j}}{h_{n}}\right)K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}1_{\left\{Y_{0}>su_{n}^{x}\right\}}\right]}{h_{n}\overline{F}^{x}(u_{n}^{x})}\leq\delta.$
@@
$\sum_{j=1}^{m}\left(m-j\right)\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}}1_{\left\{Y_{j}>su_{n}^{x}\right\}}1_{\left\{\lvert\overline{S}_{n,1}^{x}(s)\rvert>\varepsilon\sqrt{k_{n}h_{n}}\right\}}\right]$
$\leq m\sum_{j=1}^{m}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}}1_{\left\{Y_{j}>su_{n}^{x}\right\}}\right]$
$\leq\frac{m}{r_{n}}r_{n}h_{n}\overline{F}^{x}(u_{n}^{x})\sum_{j=1}^{r_{n}}\frac{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}}1_{\left\{Y_{j}>su_{n}^{x}\right\}}\right]}{h_{n}\overline{F}^{x}(u_{n}^{x})}$
$=\frac{\mathcal{O}\left({r_{n}h_{n}\overline{F}^{x}(u_{n}^{x})}\right)}{r_{n}}=o\left(r_{n}h_{n}\overline{F}^{x}(u_{n}^{x})\right).$
@@
$\sum_{i=1}^{r_{n}}\sum_{j=i+m}^{r_{n}}\mathbb{E}\left[K\left(\frac{x-X_{i}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{i}>su_{n}^{x}\right\}}1_{\left\{Y_{j}>su_{n}^{x}\right\}}1_{\left\{\lvert\overline{S}_{n,1}^{x}(s)\rvert>\varepsilon\sqrt{k_{n}h_{n}}\right\}}\right]\leq\delta r_{n}h_{n}\overline{F}^{x}(u_{n}^{x}).$
@@
$\limsup_{n\to\infty}\frac{I_{2}}{r_{n}h_{n}\overline{F}^{x}(u_{n}^{x})}\leq\delta,$
and since $\delta$ can be made arbitrarily small we conclude that $I_{2}=o(r_{n}h_{n}\overline{F}^{x}(u_{n}^{x}))$ which concludes the proof. ∎
Now we show that terms constructed from the incomplete block $\{(X_{m_{n}r_{n}+1},Y_{m_{n}r_{n}+1}),\ldots,\left(X_{n},Y_{n}\right)\}$ are asymptotically negligible.
@@
$\sup_{s\in[a,b]}\left\lvert\sum_{j=m_{n}r_{n}+1}^{n}\frac{K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}-\mathbb{E}\left[K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}\right]}{\sqrt{nh_{n}\overline{F}^{x}(u_{n}^{x})}}\right\rvert\overset{\mathrm{Pr}}{\to}0.$
@@
Since $s\mapsto K\left((x-X_{j})/h_{n}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}$ is monotonically decreasing, and $(n-m_{n}r_{n})\leq r_{n}$, i.e. the last incomplete block consists of fewer terms than a complete block, we get
$\sup_{s\in[a,b]}\left\lvert\sum_{j=m_{n}r_{n}+1}^{n}\frac{K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}-\mathbb{E}\left[K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}\right]}{\sqrt{nh_{n}\overline{F}^{x}(u_{n}^{x})}}\right\rvert$
$\leq\frac{1}{\sqrt{nh_{n}\overline{F}^{x}(u_{n}^{x})}}\Bigg\{\sum_{j=m_{n}r_{n}+1}^{n}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>au_{n}^{x}\right\}}+r_{n}\mathbb{E}\left[K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>au_{n}^{x}\right\}}\right]\Bigg\}.$
@@
$\frac{1}{nh_{n}\overline{F}^{x}(u_{n}^{x})}\operatorname{var}\left(\sum_{j=1}^{(n-m_{n}r_{n})}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>au_{n}^{x}\right\}}\right)$
$\leq\frac{1}{nh_{n}\overline{F}^{x}(u_{n}^{x})}\operatorname{var}\left(\sum_{j=1}^{r_{n}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>au_{n}^{x}\right\}}\right)=\frac{1}{m_{n}}\operatorname{var}\left(\widetilde{\mathbb{V}}^{{\dagger},x}_{n}(a)\right)\to 0,$
where we used that the joint process $\left\{\left(X_{j},Y_{j}\right)\right\}$ is stationary, Lemma 5 and Lemma 6 . Further, by condition 1,
$\frac{r_{n}}{\sqrt{nh_{n}\overline{F}^{x}(u_{n}^{x})}}\mathbb{E}\left[K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>au_{n}^{x}\right\}}\right]=\frac{r_{n}}{\sqrt{nh_{n}\overline{F}^{x}(u_{n}^{x})}}\mathcal{O}\left(h_{n}\overline{F}^{x}(u_{n}^{x})\right)=\frac{\mathcal{O}\left(r_{n}h_{n}\overline{F}^{x}(u_{n}^{x})\right)}{\sqrt{nh_{n}\overline{F}^{x}(u_{n}^{x})}}\to 0.$
@@
$\sum_{j=1}^{r_{n}}\frac{K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}}{\sqrt{nh_{n}\overline{F}^{x}(u_{n}^{x})}}.$
@@
$m_{n}\mathbb{E}\left[\left(\sum_{j=1}^{\ell_{n}}\frac{K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}}{\sqrt{nh_{n}\overline{F}^{x}(u_{n}^{x})}}\right)^{2}\right]\leq\frac{l_{n}}{r_{n}}\operatorname{var}\left(\widetilde{\mathbb{V}}^{{\dagger},x}_{n}(s)\right)\to 0.$
@@
Assume conditions 1-41-3. Then $\widetilde{\mathbb{V}}^{x}_{n}$ converges weakly to the centred Gaussian process ${\mathbb{V}}^{x}$ in $\mathbb{D}((0,\infty))$ endowed with the $J_{1}$-topology, and the process $\mathbb{V}^{x}$ admits an almost surely continuous version.
@@
$\widetilde{\mathbb{V}}^{x}_{n,1}=\sum_{\begin{subarray}{c}1\leq i\leq m_{n}\\ i\ \text{odd}\end{subarray}}\frac{\overline{S}_{n,i}^{x}}{\sqrt{nh_{n}\overline{F}^{x}(u_{n}^{x})}},\quad\ \ \widetilde{\mathbb{V}}^{x}_{n,2}=\sum_{\begin{subarray}{c}1\leq i\leq m_{n}\\ i\ \text{even}\end{subarray}}\frac{\overline{S}_{n,i}^{x}}{\sqrt{nh_{n}\overline{F}^{x}(u_{n}^{x})}}$
$\widetilde{\mathbb{V}}^{{\dagger},x}_{n,1}=\sum_{\begin{subarray}{c}1\leq i\leq m_{n}\\ i\ \text{odd}\end{subarray}}\frac{\overline{S}_{n,i}^{{\dagger},x}}{\sqrt{nh_{n}\overline{F}^{x}(u_{n}^{x})}},\quad\widetilde{\mathbb{V}}^{{\dagger},x}_{n,2}=\sum_{\begin{subarray}{c}1\leq i\leq m_{n}\\ i\ \text{even}\end{subarray}}\frac{\overline{S}_{n,i}^{{\dagger},x}}{\sqrt{nh_{n}\overline{F}^{x}(u_{n}^{x})}}$
@@
$\widetilde{\mathbb{V}}^{x,\text{incomplete}}_{n}(s):=\sum_{j=m_{n}r_{n}+1}^{n}\frac{K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}-\mathbb{E}\left[K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}\right]}{\sqrt{nh_{n}\overline{F}^{x}(u_{n}^{x})}}.$
@@
The summands of $\widetilde{\mathbb{V}}^{{\dagger},x}_{n}$ are i.i.d processes indexed on $[a,b]$ with càdlàg sample paths, the latter implying separability of each summand. Consequently we can apply Theorem 19 to show asymptotic equicontinuity. The Lindeberg condition was already shown in Theorem 10. With $v_{n}=nh_{n}\overline{F}^{x}(u_{n}^{x})$ define the random metric $\mathrm{d}_{n}$ on $[a,b]$ by
@@
$\mathbb{E}\left[\mathrm{d}_{n}^{2}(s,t)\right]=\frac{m_{n}}{nh_{n}\overline{F}^{x}(u_{n}^{x})}\mathbb{E}\left[\left\{\overline{S}_{n,i}^{{\dagger},x}(s)-\overline{S}_{n,i}^{{\dagger},x}(t)\right\}^{2}\right]$
$\leq\frac{m_{n}}{nh_{n}\overline{F}^{x}(u_{n}^{x})}\left(\mathbb{E}\left[\overline{S}_{n,i}^{{\dagger},x}(s)^{2}\right]-\mathbb{E}\left[\overline{S}_{n,i}^{{\dagger},x}(t)^{2}\right]\right)\to\left(s^{-1/\gamma(x)}-t^{-1/\gamma(x)}\right)g(x)\int K^{2}\left(u\right)\mathrm{d}u,$
@@
$\lim_{n\to\infty}\sup_{\begin{subarray}{c}s,t\in[a,b]\\ \lvert s-t\rvert<\delta_{n}\end{subarray}}\mathbb{E}\left[\mathrm{d}_{n}^{2}(s,t)\right]\leq\lim_{n\to\infty}\sup_{\begin{subarray}{c}s,t\in[a,b]\\ \lvert s-t\rvert<\delta_{n}\end{subarray}}\left\{\frac{m_{n}}{nh_{n}\overline{F}^{x}(u_{n}^{x})}\left(\mathbb{E}\left[\overline{S}_{n,i}^{{\dagger},x}(s)^{2}\right]-\mathbb{E}\left[\overline{S}_{n,i}^{{\dagger},x}(t)^{2}\right]\right)\right\}=0.$
@@
$\mathbb{Z}_{n}\left(\mathcal{E}_{s}\right)$ $=v_{n}^{-1}\sum_{i=1}^{m_{n}}\mathcal{E}_{s}\left((u_{n}^{x})^{-1}\boldsymbol{Y}^{{\dagger}}_{(i-1)r_{n}+1,ir_{n}},h_{n}^{-1}\left(\boldsymbol{x}-\boldsymbol{X}^{{\dagger}}_{(i-1)r_{n}+1,ir_{n}}\right)\right)=\frac{1}{nh_{n}\overline{F}^{x}(u_{n}^{x})}\sum_{i=1}^{m_{n}}S_{n,i}^{{\dagger},x}(s),$
@@
$\widetilde{d}_{n}^{2}\left(\mathcal{E}_{s},\mathcal{E}_{t}\right)$ $=v_{n}^{-1}\sum_{i=1}^{m_{n}}\Biggl\{\mathcal{E}_{s}\left((u_{n}^{x})^{-1}\boldsymbol{Y}^{{\dagger}}_{(i-1)r_{n}+1,ir_{n}},h_{n}^{-1}\left(\boldsymbol{x}-\boldsymbol{X}^{{\dagger}}_{(i-1)r_{n}+1,ir_{n}}\right)\right)$
$\qquad\qquad\quad-\mathcal{E}_{t}\left((u_{n}^{x})^{-1}\boldsymbol{Y}^{{\dagger}}_{(i-1)r_{n}+1,ir_{n}},h_{n}^{-1}\left(\boldsymbol{x}-\boldsymbol{X}^{{\dagger}}_{(i-1)r_{n}+1,ir_{n}}\right)\right)\Biggr\}^{2},$
@@
$\boldsymbol{G}\left(\boldsymbol{y},\boldsymbol{x}\right)=\sup_{s\in[a,b]}\left\lvert\mathcal{E}_{s}\left(\boldsymbol{y},\boldsymbol{x}\right)\right\rvert=\sup_{s\in[a,b]}\left\lvert\sum_{j\in\mathbb{Z}}1_{\left\{y_{j}>s\right\}}K(x_{j})\right\rvert=\sum_{j\in\mathbb{Z}}1_{\left\{y_{j}>a\right\}}K(x_{j}).$
@@
$\limsup_{n\to\infty}\frac{m_{n}}{nh_{n}\overline{F}^{x}(u_{n}^{x})}\mathbb{E}\left[\boldsymbol{G}^{2}\left((u_{n}^{x})^{-1}\boldsymbol{Y}^{{\dagger}}_{1,r_{n}},h_{n}^{-1}\left(\boldsymbol{x}-\boldsymbol{X}^{{\dagger}}_{1,r_{n}}\right)\right)\right]$
$=\limsup_{n\to\infty}\frac{m_{n}}{nh_{n}\overline{F}^{x}(u_{n}^{x})}\mathbb{E}\left[\left(\sum_{j=1}^{r_{n}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>au_{n}^{x}\right\}}\right)^{2}\right]$
$=\limsup_{n\to\infty}\operatorname{var}\left(\widetilde{\mathbb{V}}^{{\dagger},x}_{n}(a)\right)=a^{-1/\gamma(x)}g(x)\int K^{2}(u)\mathrm{d}u<\infty.$